OpenStudy (anonymous):

x^2-8x-y^2=-12 graph

OpenStudy (mertsj):

Complete the square : x^2-8x -y^2=-12 x^2-8x+16 -y^2=-12+16 (x-4)^2-y^2=4 Divide both sides by 4 (x-4)^2/4-y^2/4=1 This is a hyperbola whose center is (4,0)